A spring of force constant 1500N/m is acted upon by a force constant of 75N.Calculate the potential energy stored in the spring.
1
Expert's answer
2018-04-29T13:46:07-0400
First, let’s calculate the compression of the spring from the Hooke’s law:
F=kx,
here, F=75 N is the force acted on the spring, k=1500 N⁄m is the force constant of the spring, x is the compression of the spring.
Then, we get:
x=F/k=(75 N)/(1500 N/m)=0.05 m.
Finally, we can find the potential energy stored in the spring:
PE=1/2 kx^2=1/2∙1500 N/m∙(0.05 m)^2=1.875 J.